a man invested $25,000, part of it was invested at 3% and the rest at 4%. the total annual income from two investments is $950.00. how much was invested at each of these rates?
Answer by checkley77(12844)

.04X+.03(25,000-X)=950
.04X+750-.03X=950
.01X=950-750
.01X=200
X=200/.01
X=20,000 AMOUNT INVESTED @ 4%
25,000-20,000=5,000 INVESTED @ 3%
PROOF:
.04*20,000+.03*5,000=950
800+150=950
950=950